# Quaternion tensor low rank approximation

### v2

@@ 2.1 Basic quaternion algebra and notations
The quaternion skew-field $\mathbb{H}$ is an associative but non-commutative algebra of rank 4 over $\mathbb{R}$, and 2 over $\mathbb{C}$. For a quaternion $\dot{q}$, we denote $\mathfrak{R}(\dot{q}):=q_{0}$, and $\mathfrak{J}(\dot{q}):=q_{1}\mathbf{i}+q_{2}\mathbf{j}+q_{3}\mathbf{k}$ as the real and imaginary part of a quaternion, respectively. If the real part is zero, it is called, a pure quaternion. The conjugate of $\dot{q}$ is given by $\dot{\widebar{q}}=\mathfrak{R}(\dot{q})-\mathfrak{J}(\dot{q})$, and its norm is $||\dot{q}||=\sqrt{\dot{q}\dot{\widebar{q}}}=\sqrt{q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}}$.
Let (.$(.)^{T},(.)^{H}$ denote the transpose and the conjugate transpose, respectively.
@@ 2.2 Quaternion matrix and tensor
Definition 2.2.
Given the QSVD $\dot{Q}=\dot{U}S\dot{V}^{H}$, with $S=\operatorname{diag}(\sigma_{i})$. The nuclear norm of $\dot{Q}$ is $||\dot{Q}||_{*}:=\sum_{i}\sigma_{i}$, i.e, the sum of its singular values.
Definition 2.3 (Mode-k unfolding).
@@
Definition 2.4 (Tucker Rank).
@@
The corresponding Frobenius norm $||\dot{\mathcal{Q}}||_{F}:=\sqrt{\langle\dot{\mathcal{Q}},\dot{\mathcal{Q}}\rangle}=\sqrt{\sum_{i_{1},\ldots,i_{N}}||\dot{q}_{i_{1}i_{2}\ldots i_{N}}||^{2}}$.
@@ 2.3 The QT-product
Definition 2.6.
[24] Given two N-th order ($N\geq 3$) quaternion tensors $\dot{\mathcal{Q}}\in\mathbb{H}^{I_{1}\times J\times I_{3}\ldots I_{N}},\dot{\mathcal{P}}\in\mathbb{H}^{J\times I_{2}\times\ldots I_{N}}$ and invertible matrices $\dot{M}_{i}\in\mathbb{H}^{I_{i}\times I_{i}},i\in[|3,N|]$, the ×$\times_{QT}$ product is defined as
@@
with ×$\times_{QT}$ is the face-wised product defined [17] by frontal slices as
@@
  • The identity quaternion tensor ˙$\dot{\mathcal{I}}\in\mathbb{H}^{J\times J\times I_{3}\ldots\times I_{N}}$ satisfies ˙^$\widehat{\dot{\mathcal{I}}}(:,:,i_{3},\ldots,i_{N})=I_{J}$.
@@
  • For third order tensors, the quaternion Tensor Nuclear Norm (QTNN) [34] as $||\dot{\mathcal{Q}}||_{*}=\sum_{i,j}\sigma_{j}(\widehat{\dot{\mathcal{Q}}}^{(i)})=\sum_{i,j}\sigma_{j}(\mathcal{S}^{(i)}).$
@@
1:$\widehat{\dot{\mathcal{Q}}}:=L(\dot{\mathcal{Q}})=\dot{\mathcal{Q}}\times_{3}\dot{M}$$\triangleright$ Transformed domain
@@
5:$\dot{\mathcal{U}}=L^{-1}(\widehat{\dot{\mathcal{U}}}),\mathcal{S}=L^{-1}(\widehat{\mathcal{S}}),\dot{\mathcal{V}}=L^{-1}(\widehat{\dot{\mathcal{V}}})$.
@@ 2.4 Introducing the Non-Convex surrogate functions
Definition 2.10.
  • Let $\dot{Q}\in\mathbb{H}^{I_{1}\times I_{2}}$, we define the $\phi_{\gamma}$ norm as $||\dot{Q}||_{\phi_{\gamma}}:=\sum_{i}\phi_{\gamma}(\sigma_{i}(\dot{Q}))$.
  • Let $\mathcal{\dot{Q}}\in\mathbb{H}^{I_{1}\times I_{2}\times I_{3}}$, the QT-$\phi_{\gamma}$ norm is defined as, $||\mathcal{\dot{Q}}||_{\phi_{\gamma}}=\sum_{i}||\widehat{\dot{\mathcal{Q}}}^{(i)}||_{\phi_{\gamma}}=\sum_{i,j}\phi_{\gamma}(\sigma_{j}^{(i)})$.
The Non-Convex surrogate function norms has some important properties, such as, the Unitary invariant, i,e $||\dot{Q}||_{\phi_{\gamma}}=||S||_{\phi_{\gamma}}$, where $S$ is the rectangular diagonal matrix of the QSVD decomposition, as well as convergence to the rank or to the nuclear norm for specific parameters for some of these methods. The unitary invariance is also established for the QT-$\phi_{gamma}$.
@@
Proposition 2.12.
$\sigma^{(t+1)}=$ $\arg\min_{x\geq 0}\frac{1}{2}||x-\sigma^{(t)}||_{2}^{2}+\lambda\partial\phi_{\gamma}(\sigma^{(t)})x$ (5)
@@
$\arg\min_{\mathcal{\dot{X}}}\dfrac{1}{2}||\mathcal{\dot{X}}-\mathcal{\dot{Q}}||_{F}^{2}+\lambda||\mathcal{\dot{X}}||_{\phi_{\gamma}}=\dot{\mathcal{U}}\times_{QT}\mathcal{S}^{*}\times_{QT}\dot{\mathcal{V}}^{H}.$ (6)
@@
$\sum_{i}||\widehat{\dot{\mathcal{X}}}^{(i)}-\widehat{\dot{\mathcal{Q}}}^{(i)}||_{F}^{2}+\lambda||\widehat{\dot{\mathcal{X}}}^{(i)}||_{\phi_{\gamma}},$
which leads to solving $I_{3}$ independent problems; $||\widehat{\dot{\mathcal{X}}}^{(i)}-\widehat{\dot{\mathcal{Q}}}^{(i)}||_{F}^{2}+\lambda||\widehat{\dot{\mathcal{X}}}^{(i)}||_{\phi_{\gamma}}$, where the solution is given in proposition 2.12.
@@
1:$\left[\dot{\mathcal{U}},\mathcal{S},\dot{\mathcal{V}}\right]$= $(\dot{\mathcal{Q}})$.
@@
5:              Compute $\sigma_{j}^{*(i)}$ from $\sigma_{j}^{(i)}$ via 2.12.
@@
10:$\mathcal{X}=\dot{\mathcal{U}}\times_{QT}\mathcal{S}^{*}\times_{QT}\dot{\mathcal{V}}^{H}$.
Proposition 2.15 ([34]).
@@
$\mathfrak{Shrink}_{\lambda}(\dot{\mathcal{Q}})$ $:=\arg\min_{\dot{\mathcal{X}}}||\dot{\mathcal{X}}-\dot{\mathcal{Q}}||_{F}^{2}+\lambda||\dot{\mathcal{X}}||_{1}$
@@
with the |.||.|| . | is the element wise norm, and the function $\operatorname{sign}$ is an element wise operator defined as,
$\operatorname{sign}(\dot{q}_{i,j,k})=\begin{cases}\dfrac{\dot{q}_{i,j,k}}{||\dot{q}_{i,j,k}||}\text{ if }|\dot{q}_{i,j,k}|\neq 0,\\ 0,\text{ otherwise.}\end{cases}$
@@ 3 Low rank tensor completion
with the non-negatives weights $\alpha_{k}$, satisfying $\sum_{k}^{N}\alpha_{k}=1$, and the linear operation .$_{\Omega}$ keeps the entries in $\Omega$ and zeros out others.
$\min_{\dot{\mathcal{Q}}}\sum_{k}^{N}\alpha_{k}||\dot{Q}_{[k]}||_{*}$ (8)
[…]
@@ 3.1 Solution of the proposed model
[…]
@@
$=\arg\min_{\dot{Q}_{k}}\alpha_{k}||\dot{Q}_{k}||_{\phi_{\gamma}}+\mathfrak{R}\left(\left\langle\dot{F}_{k}^{(\tau)},\dot{P}_{[k]}^{(\tau+1)}-\dot{Q}_{k}\right\rangle\right)+\frac{\beta_{k}^{(\tau)}}{2}\left\|\dot{P}_{[k]}^{(\tau+1)}-\dot{Q}_{k}\right\|_{F}^{2}$
$=\arg\min_{\dot{Q}_{k}}\alpha_{k}||\dot{Q}_{k}||_{\phi_{\gamma}}+\frac{\beta_{k}^{(\tau)}}{2}\left\|\dot{P}_{[k]}^{(\tau+1)}-\dot{Q}_{k}+\frac{1}{\beta_{k}^{(\tau)}}\dot{F}_{k}^{(\tau)}\right\|_{F}^{2}$
@@ 3.2 Convergence analysis
In this part, we show the convergence is guaranteed using theorem 3.5, first we need the following two lemmas.
Lemma 3.1.
@@
$=$ $\alpha_{k}\partial_{\dot{Q}_{k}}\left(||\dot{Q}_{k}||_{\phi_{\gamma}}+\frac{\beta_{k}^{(\tau)}}{2}\left\|\dot{P}_{[k]}^{(\tau+1)}-\dot{Q}_{k}+\frac{1}{\beta_{k}^{(\tau)}}\dot{F}_{k}^{(\tau)}\right\|_{F}^{2}\right)|_{\dot{Q}_{k}^{(\tau+1)}}$
$=$ $\alpha_{k}\partial_{\dot{Q}_{k}}||\dot{Q}_{k}||_{\phi_{\gamma}}|_{\dot{Q}_{k}^{(\tau+1)}}+\beta_{k}^{(\tau)}\left(\dot{Q}_{k}^{(\tau+1)}-\dot{P}_{[k]}^{(\tau+1)}-\frac{1}{\beta_{k}^{(\tau)}}\dot{F}_{k}^{(\tau)}\right)$
$=$ $\alpha_{k}\partial_{\dot{Q}_{k}}||\dot{Q}_{k}||_{\phi_{\gamma}}|_{\dot{Q}_{k}^{(\tau+1)}}-\dot{F}_{k}^{(\tau+1)}.$
As the subgradient $\partial_{\dot{Q}_{k}}||\dot{Q}_{k}||_{\phi_{\gamma}}|_{\dot{Q}_{k}^{(\tau+1)}}$ is bounded, hence, the sequence $\left\{\dot{F}_{k}^{(\tau)}\right\}_{\tau}^{\infty},\;k=1,\ldots,N,$ is bounded.
Lemma 3.3.
@@
$\;\qquad+\sum_{k}\frac{\beta_{k}^{(\tau)}-\beta_{k}^{(\tau-1)}}{2}\left\|\dot{P}_{[k]}^{(\tau)}-\dot{Q}_{k}^{(\tau)}\right\|_{F}^{2}+\mathfrak{R}\left(\left\langle\dot{F}_{k}^{(\tau)}-\dot{F}_{k}^{(\tau-1)},\dot{P}_{[k]}^{(\tau)}-\dot{Q}_{k}^{(\tau)}\right\rangle\right)$
$\;\qquad=\mathfrak{L}\left(\dot{\mathcal{P}}^{(\tau)},\left\{\dot{Q}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(\tau-1)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(\tau-1)}\right\}_{k=1}^{N}\right)+\sum_{k}^{N}\frac{\beta_{k}^{(\tau)}+\beta_{k}^{(\tau-1)}}{2\beta_{k}^{(\tau-1)^{2}}}\left\|\dot{F}_{k}^{(\tau)}-\dot{F}_{k}^{(\tau-1)}\right\|_{F}^{2}.$
@@
$\;\qquad\leq\mathfrak{L}\left(\dot{\mathcal{P}}^{(\tau)},\left\{\dot{Q}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(\tau)}\right\}_{k=1}^{N}\right)$
$\;\qquad\leq\mathfrak{L}\left(\dot{\mathcal{P}}^{(\tau)},\left\{\dot{Q}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(\tau-1)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(\tau-1)}\right\}_{k=1}^{N}\right)+\sum_{k}^{N}\frac{\beta_{k}^{(\tau)}+\beta_{k}^{(\tau-1)}}{2\beta_{k}^{(\tau-1)^{2}}}\left\|\dot{F}_{k}^{(\tau)}-\dot{F}_{k}^{(\tau-1)}\right\|_{F}^{2}$
$\;\qquad\leq\mathfrak{L}\left(\dot{\mathcal{P}}^{(1)},\left\{\dot{Q}_{k}^{(1)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(0)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(0)}\right\}_{k=1}^{N}\right)+\sum_{k}^{N}\sum_{t=1}^{\tau}\frac{\beta_{k}^{(t)}+\beta_{k}^{(t-1)}}{2\beta_{k}^{(t-1)^{2}}}\left\|\dot{F}_{k}^{(t)}-\dot{F}_{k}^{(t-1)}\right\|_{F}^{2},$
@@
Using Lemma 3.3, along with the $\beta$-Convergence condition, then, the right-hand side is upper bounded. Since each term of the Lagrangian is non-negative, hence, we can deduce the result.
@@ 4 Low rank tensor problem
$\min_{\dot{\mathcal{Q}},\dot{\mathcal{S}}}||\dot{\mathcal{Q}}||_{\phi_{\gamma}}+\lambda||\dot{\mathcal{S}}||_{1}$ (20)
@@ 4.1 Solution of the proposed model
$\mathfrak{L}\left(\dot{\mathcal{Q}},\dot{\mathcal{S}},\dot{\mathcal{Y}}\right)=||\dot{\mathcal{Q}}||_{\phi_{\gamma}}+\lambda||\dot{\mathcal{S}}||_{1}+\mathfrak{R}\left(\left\langle\mathcal{\dot{Y}},\mathcal{\dot{X}}-\mathcal{\dot{Q}}-\mathcal{\dot{S}}\right\rangle\right)+\frac{\beta}{2}\left\|\mathcal{\dot{X}}-\mathcal{\dot{Q}}-\mathcal{\dot{S}}\right\|_{F}^{2},$
@@
$\dot{\mathcal{Q}}^{(\tau+1)}$ $=\arg\min_{\dot{\mathcal{Q}}}||\dot{\mathcal{Q}}||_{\phi_{\gamma}}+\frac{\beta^{(\tau)}}{2}\left\|\mathcal{\dot{Q}}+\mathcal{\dot{S}}^{(\tau)}-\mathcal{\dot{X}}-\dfrac{\mathcal{\dot{Y}}^{(\tau)}}{\beta^{(\tau)}}\right\|_{F}^{2}$ (25)
@@
$\dot{\mathcal{S}}^{(\tau+1)}$ $=\arg\min_{\dot{\mathcal{S}}}\lambda||\dot{\mathcal{S}}||_{1}+\frac{\beta^{(\tau)}}{2}\left\|\mathcal{\dot{S}}+\mathcal{\dot{Q}}^{(\tau+1)}-\mathcal{\dot{X}}-\frac{\mathcal{\dot{Y}}^{(\tau)}}{\beta^{(\tau)}}\right\|_{F}^{2}$ (26)

### v3

# Quaternion tensor low-rank approximation using a family of non-convex norms

@@ 2.1 Basic quaternion algebra and notations
The quaternion skew-field $\mathbb{H}$ is an associative but non-commutative algebra of rank 4 over $\mathbb{R}$, and 2 over $\mathbb{C}$. For a quaternion $\dot{q}$, we denote $\mathfrak{R}(\dot{q}):=q_{0}$, and $\mathfrak{J}(\dot{q}):=q_{1}\mathbf{i}+q_{2}\mathbf{j}+q_{3}\mathbf{k}$ as the real and imaginary part of a quaternion, respectively. If the real part is zero, it is called, a pure quaternion. The conjugate of $\dot{q}$ is given by $\dot{\widebar{q}}=\mathfrak{R}(\dot{q})-\mathfrak{J}(\dot{q})$, and its norm is $\dot{q}||=\sqrt{\dot{q}\dot{\widebar{q}}}=\sqrt{q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}}$.
Let $(.)^{T},(.)^{H}$ denote the transpose and the conjugate transpose, respectively.
@@ 2.2 Quaternion matrix and tensor
Definition 2.
Given the QSVD $\dot{Q}=\dot{U}S\dot{V}^{H}$, with $S=\operatorname{diag}(\sigma_{i})$. The nuclear norm of $\dot{Q}$ is $\dot{Q}||_{*}:=\sum_{i}\sigma_{i}$, i.e, the sum of its singular values.
Definition 3 (Mode-k unfolding).
@@
Definition 4 (Tucker Rank).
@@
[…]
@@ 2.3 The QT-product
Definition 6.
[24] Given two N-th order ($N\geq 3$) quaternion tensors $\dot{\mathcal{Q}}\in\mathbb{H}^{I_{1}\times J\times I_{3}\ldots I_{N}},\dot{\mathcal{P}}\in\mathbb{H}^{J\times I_{2}\times\ldots I_{N}}$ and invertible matrices $\dot{M}_{i}\in\mathbb{H}^{I_{i}\times I_{i}},i\in[|3,N|]$, the $\times_{QT}$ product is defined as
@@
with $\times_{QT}$ is the face-wised product defined [17] by frontal slices as
@@
  • The identity quaternion tensor $\dot{\mathcal{I}}\in\mathbb{H}^{J\times J\times I_{3}\ldots\times I_{N}}$ satisfies $\widehat{\dot{\mathcal{I}}}(:,:,i_{3},\ldots,i_{N})=I_{J}$.
@@
  • For third order tensors, the quaternion Tensor Nuclear Norm (QTNN) [34] as $\dot{\mathcal{Q}}||_{*}=\sum_{i,j}\sigma_{j}(\widehat{\dot{\mathcal{Q}}}^{(i)})=\sum_{i,j}\sigma_{j}(\mathcal{S}^{(i)}).$
@@
[…]
@@
[…]
@@ 2.4 Introducing the Non-Convex surrogate functions
Definition 10.
  • Let $\dot{Q}\in\mathbb{H}^{I_{1}\times I_{2}}$, we define the $\phi_{\gamma}$ norm as $\dot{Q}||_{\phi_{\gamma}}:=\sum_{i}\phi_{\gamma}(\sigma_{i}(\dot{Q}))$.
[…]
@@
A common method to solve the problem (minimize $f(x)+g(x)$) with Non-Convex regularizer, is Difference of Convex functions (DC) [11], it minimizes the Non-Convex function g(.)-g(.), based on the assumption that both the functions (f(.),g(.)(f(.),-g(.) are convex. In each iteration, DC programming linearizes g(.)-g(.) at $x=x^{t}$, and minimizes a relaxed function.
Proposition 12.
@@
[…]
@@
[…]
@@
[…]
@@
5:     Compute $\sigma_{j}^{*(i)}$ from $\sigma_{j}^{(i)}$ via 12.
@@
[…]
Proposition 14 ([34]).
@@
[…]
@@
with the $|.|$ is the element wise norm, and the function $\operatorname{sign}$ is an element wise operator defined as,
[…]
@@ 3 Low rank tensor completion
$\text{s.t }\dot{\mathcal{Q}}_{\Omega}=\dot{\mathcal{O}}_{\Omega},$
with the non-negatives weights $\alpha_{k}$, satisfying $\sum_{k}^{N}\alpha_{k}=1$, and the linear operation $_{\Omega}$ keeps the entries in $\Omega$ and zeros out others.
@@
$\text{s.t }\dot{\mathcal{Q}}_{\Omega}=\dot{\mathcal{O}}_{\Omega},$
@@ 3.1 Solution of the proposed model
$\text{s.t }\dot{P}_{[k]}=\dot{Q}_{k},k=1,\ldots,N,\;\dot{\mathcal{P}}_{\Omega}=\dot{\mathcal{O}}_{\Omega}.$
@@
[…]
@@ 3.2 Convergence analysis
In this part, we show the convergence is guaranteed using theorem 17, first we need the following two lemmas.
Lemma 15.
@@
[…]
Lemma 16.
@@
$\;\hskip 18.49988pt+\sum_{k}\frac{\beta_{k}^{(\tau)}-\beta_{k}^{(\tau-1)}}{2}\left\|\dot{P}_{[k]}^{(\tau)}-\dot{Q}_{k}^{(\tau)}\right\|_{F}^{2}+\mathfrak{R}\left(\left\langle\dot{F}_{k}^{(\tau)}-\dot{F}_{k}^{(\tau-1)},\dot{P}_{[k]}^{(\tau)}-\dot{Q}_{k}^{(\tau)}\right\rangle\right)$
$\;\hskip 18.49988pt=\mathfrak{L}\left(\dot{\mathcal{P}}^{(\tau)},\left\{\dot{Q}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(\tau-1)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(\tau-1)}\right\}_{k=1}^{N}\right)+\sum_{k}^{N}\frac{\beta_{k}^{(\tau)}+\beta_{k}^{(\tau-1)}}{2\beta_{k}^{(\tau-1)^{2}}}\left\|\dot{F}_{k}^{(\tau)}-\dot{F}_{k}^{(\tau-1)}\right\|_{F}^{2}.$
@@
$\;\hskip 18.49988pt\leq\mathfrak{L}\left(\dot{\mathcal{P}}^{(\tau)},\left\{\dot{Q}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(\tau)}\right\}_{k=1}^{N}\right)$
$\;\hskip 18.49988pt\leq\mathfrak{L}\left(\dot{\mathcal{P}}^{(\tau)},\left\{\dot{Q}_{k}^{(\tau)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(\tau-1)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(\tau-1)}\right\}_{k=1}^{N}\right)+\sum_{k}^{N}\frac{\beta_{k}^{(\tau)}+\beta_{k}^{(\tau-1)}}{2\beta_{k}^{(\tau-1)^{2}}}\left\|\dot{F}_{k}^{(\tau)}-\dot{F}_{k}^{(\tau-1)}\right\|_{F}^{2}$
$\;\hskip 18.49988pt\leq\mathfrak{L}\left(\dot{\mathcal{P}}^{(1)},\left\{\dot{Q}_{k}^{(1)}\right\}_{k=1}^{N},\left\{\dot{F}_{k}^{(0)}\right\}_{k=1}^{N},\left\{\beta_{k}^{(0)}\right\}_{k=1}^{N}\right)+\sum_{k}^{N}\sum_{t=1}^{\tau}\frac{\beta_{k}^{(t)}+\beta_{k}^{(t-1)}}{2\beta_{k}^{(t-1)^{2}}}\left\|\dot{F}_{k}^{(t)}-\dot{F}_{k}^{(t-1)}\right\|_{F}^{2},$
@@
Using Lemma 16, along with the $\beta$-Convergence condition, then, the right-hand side is upper bounded. Since each term of the Lagrangian is non-negative, hence, we can deduce the result. ∎
@@ 4.1 Solution of the proposed model
[…]
@@
[…]
@@
[…]